Given that log2 0.3010,log3 0.4771 . find log10^75 + log0.1^64
Answer by Alan3354(69443)

Given that log2 0.3010,log3 0.4771 . find log10^75 + log0.1^64
------------
Log(10^75) = 75
---
log((10^-1)^64) = log(10^(-64)) = -64
=======
75 - 64 = 11